A  is placed in a cylinder and has a pressure of 33.3 kPa, a temperature of 25 oC, and a volume of 0.345 L. The pressure of the gas is increased to 40 kPa, and the volume is increased to 0.50 L. What is the temperature of the gas?


1
Expert's answer
2021-07-20T11:29:01-0400

"P_1=33.3kPa"

"T_1=273+25=298k"

"V_1=0.345L"



"P_2=40kPa"

"T_2=?"

"V_2=0.50L"


"\frac{P_1V_1}{T_1}
=\frac{P_2V_2}{T_2}"


Hence


"T_2=\frac{P_2V_2T_1}{P_1V_1}"


"T_2=\frac{40×0.50×298}{33.3×0.345}=518.78K"